using A=P(1+r/n)^nt how long must $4700 be in a bank at 5% compounded annually to become $6944.04?
Answer by nerdybill(7384)

using A=P(1+r/n)^nt how long must $4700 be in a bank at 5% compounded annually to become $6944.04?
.
If compounded annually, n would be 1 so now we have
A=P(1+r)^t
.
A = 6944.04
P = 4700
r = .05
.
A=P(1+r)^t
6944.04 = 4700(1+.05)^t
6944.04 = 4700(1.05)^t
6944.04 = 4700(1.05)^t
1.4775 = (1.05)^t
Taking log base 1.05 of each side AND applying the "change of base":
log(1.4775)/log(1.05) = t
8 years = t